OpenStudy (anonymous):

0.2t=0.4+0.7t what is soulution set?

OpenStudy (saifoo.khan):

add like terms

OpenStudy (saifoo.khan):

0.7t - 0.2t = 0.4

OpenStudy (saifoo.khan):

0.5t = 0.4 t = 0.4/0.5

OpenStudy (saifoo.khan):

therefore: t = 4/5

OpenStudy (anonymous):

is it negative or positive when you take the 0.7 from 0.2?

OpenStudy (saifoo.khan):

Oh, wait. sorry!

OpenStudy (saifoo.khan):

t = -4/5

OpenStudy (anonymous):

lol..lol thank you can you help me with I7x-4/2x-3I=4 ?

OpenStudy (saifoo.khan):

yes i can. :0

OpenStudy (saifoo.khan):

:)

OpenStudy (saifoo.khan):

Please post it again on left. i will come over there!